Function Value

Please help me solve this problem

Assume that f(3)= -8 (negative 8)

Find another function value, in the form F(a)= b, if:

a) the graph of y= f(x) is symmetric with respect to the y-axis

b) f( -x )= -f (x) for all x
Answer #1

f(3) = -8 WJ If f()= Summetric with 4-axis then f (oc)= f(x) & f(-x)=of (oc) f(3) = -8 f(-3) = -8 6] F (-x) = -f(oc) then f(3